# Sampling and estimation on manifolds using the Langevin diffusion

### v2

@@ 4.5. Finite-time convergence
Although this paper is focused on the use of SDEs to sample from a given distribution on $M$, and hence on the estimators for ergodic limits from Section 4, for some applications it may also be of interest to discuss weak convergence of a Riemannian Langevin algorithm at a finite time $T$; this, for example, can be used to evaluate quantities related to transient behaviour of diffusions on manifold; solving linear parabolic PDEs on manifolds by exploiting probabilistic representations (see (2.15)). It is straightforward to prove the following weak-sense convergence theorem along the lines of the proof of Theorem 4.1.
@@
(4.14) $dX=a(t,X)dt+dB^{M}(t),\quad X(t_{0})=x,$
@@ 5. The case of non-compact manifolds
The condition (2.9) is sufficient to ensure stochastic completeness of a diffusion with generator $\mathcal{A}$ in (2.6); recall that it is automatically satisfied for compact $M$. Hence here we state (2.9) as the below assumption.
@@
Ergodicity, on the other hand, requires more. The Bakry–Èmery criterion (Bakry and Émery, 1985) demands the existence of a $\kappa>0$ such that
@@
which jointly controls both curvature of $M$ and drift of the Brownian motion, and is more restrictive than (2.9). For example, when $M=\mathbb{R}^{q}$, the criterion (5.1) reduces to $\mathrm{Hess}^{\phi}\geq 2\kappa g$ for $\kappa>0$, and implies that the density $p$ is $2\kappa$-strongly log-concave. The Bakry–Èmery criterion provides a sufficient condition for exponential ergodicity of the diffusion (2.7) on non-compact $M$, but is redundant for compact $M$. Recall that by construction, $\mu_{\phi}$ is the invariant measure of the diffusion if it is ergodic.
A few intermediate conditions signpost the route to exponential ergodicity of the Langevin diffusion that satisfies the Bakry–Èmery criterion. The criterion ensures that the semigroup $\{P_{t}\}$ of operators corresponding to the diffusion satisfy the Log-Sobolev inequality with constant $2\kappa$ (Wang, 2009), which then implies that they satisfy the Poincaré inequality: $\mathrm{Var}_{\mu_{\phi}}(\varphi)\leq C\mathcal{E}(\varphi,\varphi)$ for a Dirichlet form $\mathcal{E}$ and constant $C$. As a consequence, one gets $|{P_{t}}|_{\text{op}}\leq e^{-t/C}$, which implies exponential ergodicity.
The Bakry–Èmery criterion, however, is too stringent for our needs, as it precludes the possibility of sampling from a sizeable class of interesting probability measures on $M$. It is instead more natural to impose separate conditions on $\mathrm{Ric}$ and $\mathrm{Hess}^{\phi}$. The following assumption is sufficient for the semigroup of operators corresponding to the Langevin diffusion to satisfy the Log-Sobolev inequality (Wang, 2009, Theorem 1.1).
@@
which together with constant diffusion coefficient is sufficient for exponential ergodicity of the corresponding SDE in $\mathbb{R}^{q}$ (Khasminskii, 2012; Roberts and Tweedie, 1996).
@@ 5.1. Examples
Let $M$ be the Poincarè model $\mathbb{H}^{q}$ for the $q$-dimensional hyperbolic space with metric $g=\frac{1}{x_{q}^{2}}\sum_{i=1}^{q}\text{d}x_{i}^{2}$, under which $M$ has negative sectional curvature equal to $-1$ everywhere. From the Cartan–Hadamard theorem, it is diffeomorphic to $\mathbb{R}^{q}$. Consider the Riemannian–Gaussian distribution with density $p(x)\propto\exp(-\frac{\rho(x,o)^{2}}{2\sigma^{2}})$, with respect to $\mathrm{dvol}_{g}$, where $o\in M$ is a fixed point and $\sigma>0$; for brevity, let $\rho:=\rho(x,o)$ be the distance function to the fixed $o$.
The Hessian of the squared distance satisfies $\mathrm{Hess}^{\rho^{2}}\geq 2g$ (Greene and Wu, 2006), and since $\mathbb{H}^{q}$ is an Einstein manifold, its Ricci curvature $\mathrm{Ric}=-(q-1)g$ has a lower bound. Assumption Assumption is satisfied since $\mathrm{Ric}+\mathrm{Hess}^{\phi}\geq(\frac{1}{\sigma^{2}}-q+1)g$, ensuring stochastic completeness. Note that the lower bound need not be positive as $\sigma>0$ is unbounded, therefore the Bakry–Èmery criterion may be violated for some values of $\sigma$.
The lower bound on the Hessian means that we can restrict attention to a geodesic ball centred at $o$ of finite radius to choose the values $c=(q-1)$, $b=((1+\sqrt{2})\sigma^{2}\sqrt{q-1})^{-1/2}$ and $\delta=\frac{1}{\sigma^{2}}$ such that Assumption Assumption is satisfied in tandem. Convergence exponentially fast to $p\thinspace\mathrm{dvol}_{g}$ is thus assured for any $\sigma>0$, including for values that violate the Bakry–Èmery criterion (5.1).
@@
    Let $\phi(X)=\frac{1}{2\sigma^{2}}\rho(X,O)^{2}$ in (2.3) for $\sigma>0$ and fixed $O\in\mathcal{P}_{m}$. This choice of $\phi$ results in the Riemannian–Gaussian distribution (Section 6.2.1) on $\mathcal{P}_{m}$. It can be verified that $\mathrm{Ric}\geq-\frac{m}{4}g$, and by the Hessian comparison theorem $\mathrm{Hess}^{\rho^{2}}\geq 2g$ (e.g., Lewis, 2023, Section 4.4). For $\kappa=\left|\frac{1}{\sigma^{2}}-\frac{m}{4}\right|$, note that
@@
    and Assumption Assumption is satisfied. For (5.2) in Assumption Assumption, we use the lower bound on the Ricci curvature with $c=\frac{m}{4}$. The value of $b$ can be determined upon first calculating $\delta$ in (5.3). Since $\mathrm{Hess}^{\phi}\geq\frac{1}{\sigma^{2}}g$, we select $\delta=\frac{1}{\sigma^{2}}$ resulting in an interval $0<b<((1+\sqrt{2})\sigma^{2}\sqrt{m-1})^{-1/2}$ of possible values. Hence, Assumption Assumption is satisfied. Note that the Bakry–Èmery criterion (5.1) is only satisfied for $\sigma<\sqrt{\frac{4}{m}}$.
@@
    $\mathrm{Hess}^{f(\rho^{2})}=f^{\prime\prime}(\rho^{2})\text{d}\rho^{2}\otimes\text{d}\rho^{2}+f^{\prime}(\rho^{2})\mathrm{Hess}^{\rho^{2}}.$
@@
    $\mathrm{Hess}^{\rho^{4}}=2\text{d}\rho^{2}\otimes\text{d}\rho^{2}+2\rho^{2}\mathrm{Hess}^{\rho^{2}},$
@@
    $\mathrm{Hess}^{\phi}=D^{2}\rho^{4}-D^{2}\rho^{2}=8\rho^{2}\text{d}\rho\otimes\text{d}\rho+(2\rho^{2}-1)\mathrm{Hess}^{\rho^{2}}$
    upon using the fact that $\text{d}\rho^{2}=2\rho\text{d}\rho$. Since $\mathrm{Hess}^{\rho^{2}}\geq 2g$ and $\text{d}\rho\otimes\text{d}\rho$ is a non-negative tensor:
    $\mathrm{Hess}^{\phi}\geq 2(2\rho^{2}-1).$
    Thus, $\mathrm{Ric}+\mathrm{Hess}^{\phi}\geq\big{(}2(2\rho^{2}-1)-\frac{m}{4}\big{)}g$ and Assumption Assumption is satisfied for $\kappa=\frac{m+8}{4}$. Evidently, the Hessian is negative when $\rho<\frac{1}{\sqrt{2}}$, hence for the compact set we pick the closure of geodesic ball $\mathcal{B}_{o}(1)$ of radius $1$ and centred at $o$. Outside of this ball, the Hessian is strictly positive, and choosing $\delta=2$ ensures that (5.3) in Assumption Assumption is satisfied. This implies that choosing $b$ satisfying $0<b<(2(1+\sqrt{2})\sqrt{m-1})^{-1/2}$ takes care of Assumption Assumption. Non-convexity of $\phi$ can now be gleaned from the fact that inside of $\mathcal{B}_{o}(1/\sqrt{2})$ the Hessian is negative. To see this, we look at the Hessian contracted with vector fields $v$ such that $g(v,\nabla\rho)=0$. Then
@@ 5.2. Computational aspects
However, distributions $\phi$ and manifolds $(M,g)$ of practical interest are usually so that the globally Lipschitz assumption is violated. It is even so for Example 5.2(i) of the Riemannian–Gaussian distribution on $\mathcal{P}_{m}$.
@@
In Section 6.2 we provide numerical evidence of how the rejection technique can profitably be used by choosing a large enough closed geodesic ball, as in Assumption Assumption, on $\mathcal{P}_{m}$ using the proposed Riemannian Langevin Monte Carlo algorithm. In particular, no trajectory was rejected for $R=2.7$ in the case of the Riemannian–Gaussian distribution (Section 6.2.1) and $R=2$ for the non-convex potential (Section 6.2.2).
@@
Alternatives to the explicit Euler, that result in bounded moments, have been studied in depth for SDEs in $\mathbb{R}^{q}$ with non-globally Lipschitz coefficients. Examples include implicit methods (and in particular, the implicit Euler scheme) and explicit methods of tamed and balanced type (e.g., Mattingly et al., 2002; Hutzenthaler and Jentzen, 2015; Tretyakov and Zhang, 2013; Milstein and Tretyakov, 2021). Extensions of such methods to the non-compact manifold setting are possible and will be taken up elsewhere.
@@ 6. Numerical illustrations
We carry out numerical experiments to verify the theoretical results from Section 4 for compact manifolds $M$, and demonstrate utility of the proposed algorithm for both compact and non-compact $M$. Section 6.1 considers sampling using retractions from the von-Mises distribution on the two-dimensional unit sphere $\mathbb{S}^{2}$; sampling algorithms using extrinsic embedding coordinates was proposed by Wood (1994), and a method using intrinsic coordinates was proposed by Best and Fisher (1979). Both algorithms use rejection sampling, with the latter method appropriate only for very small or large concentrations.
Section 6.2 considers two distributions on the non-compact manifold of symmetric positive definite matrices: the Riemannian–Gaussian distribution with convex $\phi$, and a distribution with non-convex potential. Versatility of the proposed algorithm to handle both convex and non-convex potential is demonstrated.
We focus on the ensemble-averaging estimator $\hat{\mu}_{\phi,N}(\varphi)$ defined in (4.1). To this end, we simulate a large number $L$ of trajectories for a sufficiently long time $N$, and then verify that it enjoys first order of convergence in terms of time step $h$ by comparing with the exact value of $\mu_{\phi}(\varphi)$ (obtained via numerical integration). Two types of errors are reported: the estimation error
@@ 6.1. Retraction-based sampling from the von-Mises Fisher distribution on $\mathbb{S}^{2}$
which is obtained by the pulling back the Euclidean metric from $\mathbb{R}^{3}$. The inverse metric tensor assumes the form
@@
The centred von-Mises Fisher distribution $\mu_{\phi}$ on $\mathbb{S}^{2}$ is the distribution
@@
so that $\phi((r,\theta)^{\intercal})=\lambda\cos r$; performance of the algorithm for a non-centred version can be examined with obvious modifications. It was verified in Example 2.1 that the Langevin diffusion (2.7) with $\phi((r,\theta)^{\intercal})=-\lambda\cos r$ is ergodic, and converges exponentially fast to its invariant measure $\mu_{\phi}$.
@@
Since an exponential map in closed form is not available for this example, we use the retraction based on solving the above ODEs by a single step of size $\sqrt{h}$ of the standard $(1/6,1/3,1/3,1/6)$ 4th-order Runge–Kutta method (RK4) as explained in Section 3.2.
@@
We take $\varphi(x)=\sin r$ and compute, for $\lambda=1$, the exact value $\mu_{\phi}(\varphi)=\sqrt{2}\frac{I_{1}(1)\Gamma(3/2)}{I_{1/2}(1)}$, where $I_{n}(\lambda)$ is the modified Bessel function of the first kind (Lewis, 2023, Section 4.1). The algorithm was terminated at time $T=5$, sufficient to ensure that the error $|\mathbb{E}(\varphi(X(T)))$ $-\mu_{\phi}(\varphi)|$ (i.e., the error due to proximity of the SDE’s solution to the ergodic limit) is negligibly small. From Table 1 and the black line in Figure 1, we observe the first order convergence of Algorithm 6.1 as expected.
Table 1. Estimation and Monte Carlo errors for sampling from a von-Mises Fisher distribution. The parameters are $\lambda=1$, $X_{0}=(r_{0},\theta_{0})^{\intercal}=(\pi/4,\pi/4)^{\intercal}$ and $T=5$.

### v3

@@ 4.5. Finite-time convergence
Although this paper is focused on the use of SDEs to sample from a given distribution on $M$, and hence on the estimators for ergodic limits from Section 4, for some applications it may also be of interest to discuss weak convergence of a Riemannian Langevin algorithm at a finite time $T$; this, for example, can be used to evaluate quantities related to transient behavior of diffusions on manifold and solving linear parabolic PDEs on manifolds by exploiting probabilistic representations (see (2.15)). It is straightforward to prove the following weak-sense convergence theorem along the lines of the proof of Theorem 4.1.
@@
(4.14) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}X(t)=a(t,X(t))\mathrm{d}t+\mathrm{d}B^{M}(t)},\quad X(t_{0})=x,$
@@
Remark 4.3.
Results in this section demonstrate that the main techniques in the Euclidean setting on weak approximations and sampling carry over to the setting of compact manifolds $M$. In contrast, the situation with mean-square (strong) convergence is more delicate. For manifolds $M$ where every point has an empty cut-locus, the Riemannian Langevin algorithm (3.3) with $\{\xi_{n}\}$ being independent standard Gaussian random variables has mean-square (strong) order $1/2$ (Bharath et al., 2025). Although this matches the result for Euclidean SDEs with multiplicative noise (e.g. Milstein and Tretyakov, 2021, Chapter 1), the proof in the manifold setting is substantially more complex. If a manifold contains points with non-empty cut-loci (e.g. positively curved manifolds) then we believe that mean-square schemes should be based on space-time random walks; see Milstein and Tretyakov (2021, Chapter 6), and also Milstein and Tretyakov (1999) for a mean-square approximation of Euclidean SDEs in bounded domains. This is in contrast to the weak-sense convergence, where the Riemannian Langevin algorithm on compact $M$ has weak order $1$, and is agnostic to the presence or absence of cut points.
@@ 5. The Case of Non-compact Manifolds
The condition (2.9) is sufficient to ensure stochastic completeness of a diffusion with generator $\mathcal{A}$ in (2.6) (Li, 1992); recall that it is automatically satisfied for compact $M$. Hence here we state (2.9) as the below assumption.
@@
Ergodicity, on the other hand, requires more. The Bakry-Èmery criterion (Bakry and Émery, 1985) demands the existence of a $\kappa>0$ such that
@@
which jointly controls both curvature of $M$ and drift of the Brownian motion, and is more restrictive than (2.9). For example, when $M=\mathbb{R}^{q}$, the criterion (5.1) reduces to $\mathrm{Hess}^{\phi}\geq 2\kappa g$ for $\kappa>0$, and implies that the density $p$ is $2\kappa$-strongly log-concave. The Bakry-Èmery criterion provides a sufficient condition for exponential ergodicity of the diffusion (2.7) on non-compact $M$, but is redundant for compact $M$. Recall that by construction, $\mu_{\phi}$ is the invariant measure of the diffusion if it is ergodic.
A few intermediate conditions signpost the route to exponential ergodicity of the Langevin diffusion that satisfies the Bakry-Èmery criterion. The criterion ensures that the semigroup $\{P_{t}\}$ of operators corresponding to the diffusion satisfy the Log-Sobolev inequality with constant $2\kappa$ (Wang, 2009), which then implies that they satisfy the Poincaré inequality: $\mathrm{Var}_{\mu_{\phi}}(\varphi)\leq C\mathcal{E}(\varphi,\varphi)$ for a Dirichlet form $\mathcal{E}$ and constant $C$. As a consequence, one gets $|{P_{t}}|_{\text{op}}\leq e^{-t/C}$, which implies exponential ergodicity.
The Bakry-Èmery criterion, however, is too stringent for our needs, as it precludes the possibility of sampling from a sizeable class of interesting probability measures on $M$. It is instead more natural to impose separate conditions on $\mathrm{Ric}$ and $\mathrm{Hess}^{\phi}$. The following assumption is sufficient for the semigroup of operators corresponding to the Langevin diffusion to satisfy the Log-Sobolev inequality (Wang, 2009, Theorem 1.1).
@@
which together with constant diffusion coefficient is sufficient for exponential ergodicity of the corresponding SDE in $\mathbb{R}^{q}$ (where $\mathrm{Ric}=0$) (Khasminskii, 2012; Roberts and Tweedie, 1996).
@@ 5.1. Examples
Let $M$ be the Poincaré half plane model $\mathbb{H}^{q}$ for the $q$-dimensional hyperbolic space with metric $g=\frac{1}{x_{q}^{2}}\sum_{i=1}^{q}\mathrm{d}x_{i}^{2}$, under which $M$ has negative sectional curvature equal to $-1$ everywhere. From the Cartan-Hadamard theorem, it is diffeomorphic to $\mathbb{R}^{q}$. Consider the Riemannian-Gaussian distribution with density $p(x)\propto\exp(-\frac{\rho(x,o)^{2}}{2\sigma^{2}})$, with respect to $\mathrm{dvol}_{g}$, where $o\in M$ is a fixed point and $\sigma>0$; for brevity, let $\rho:=\rho(x,o)$ be the distance function to the fixed $o$.
The Hessian of the squared distance satisfies $\mathrm{Hess}^{\rho^{2}}\geq 2g$ (Greene and Wu, 2006), and since $\mathbb{H}^{q}$ is an Einstein manifold, its Ricci curvature $\mathrm{Ric}=-(q-1)g$ has a lower bound. Assumption Assumption is satisfied since $\mathrm{Ric}+\mathrm{Hess}^{\phi}\geq(\frac{1}{\sigma^{2}}-q+1)g$, ensuring stochastic completeness. Note that the lower bound need not be positive as $\sigma>0$ is unbounded, therefore the Bakry-Èmery criterion may be violated for some values of $\sigma$.
The lower bound on the Hessian means that we can restrict attention to a geodesic ball cent ered at $o$ of finite radius to choose the values $c=(q-1)$, $b=((1+\sqrt{2})\sigma^{2}\sqrt{q-1})^{-1/2}$ and $\delta=\frac{1}{\sigma^{2}}$ such that Assumption Assumption is satisfied in tandem. Convergence exponentially fast to $p\thinspace\mathrm{dvol}_{g}$ is thus assured for any $\sigma>0$, including for values that violate the Bakry-Èmery criterion (5.1).
@@
    Let $\phi(X)=\frac{1}{2\sigma^{2}}\rho(X,O)^{2}$ in (2.3) for $\sigma>0$ and fixed $O\in\mathcal{P}_{m}$. This choice of $\phi$ results in the Riemannian-Gaussian distribution (Section 6.2.1) on $\mathcal{P}_{m}$. It can be verified that $\mathrm{Ric}\geq-\frac{m}{4}g$, and by the Hessian comparison theorem $\mathrm{Hess}^{\rho^{2}}\geq 2g$ (e.g., Lewis, 2023, Section 4.4). For $\kappa=\left|\frac{1}{\sigma^{2}}-\frac{m}{4}\right|$, note that
@@
    and Assumption Assumption is satisfied. For (5.2) in Assumption Assumption, we use the lower bound on the Ricci curvature with $c=\frac{m}{4}$. The value of $b$ can be determined upon first calculating $\delta$ in (5.3). Since $\mathrm{Hess}^{\phi}\geq\frac{1}{\sigma^{2}}g$, we select $\delta=\frac{1}{\sigma^{2}}$ resulting in an interval $0<b<((1+\sqrt{2})\sigma^{2}\sqrt{m-1})^{-1/2}$ of possible values. Hence, Assumption Assumption is satisfied. Note that the Bakry-Èmery criterion (5.1) is only satisfied for $\sigma<\sqrt{\frac{4}{m}}$.
@@
    $\mathrm{Hess}^{f(\rho^{2})}=f^{\prime\prime}(\rho^{2})\mathrm{d}\rho^{2}\otimes\mathrm{d}\rho^{2}+f^{\prime}(\rho^{2})\mathrm{Hess}^{\rho^{2}}.$
@@
    $\mathrm{Hess}^{\rho^{4}}=2\mathrm{d}\rho^{2}\otimes\mathrm{d}\rho^{2}+2\rho^{2}\mathrm{Hess}^{\rho^{2}},$
@@
    $\mathrm{Hess}^{\phi}=D^{2}\rho^{4}-D^{2}\rho^{2}=8\rho^{2}\mathrm{d}\rho\otimes\mathrm{d}\rho+(2\rho^{2}-1)\mathrm{Hess}^{\rho^{2}}$
    upon using the fact that $\mathrm{d}\rho^{2}=2\rho\mathrm{d}\rho$. Since $\mathrm{Hess}^{\rho^{2}}\geq 2g$ and $\mathrm{d}\rho\otimes\mathrm{d}\rho$ is a non-negative tensor:
    $\mathrm{Hess}^{\phi}\geq 2(2\rho^{2}-1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g}.$
    Thus, $\mathrm{Ric}+\mathrm{Hess}^{\phi}\geq\big{(}2(2\rho^{2}-1)-\frac{m}{4}\big{)}g$ and Assumption Assumption is satisfied for $\kappa=\frac{m+8}{4}$. Evidently, the Hessian is negative when $\rho<\frac{1}{\sqrt{2}}$, hence for the compact set we pick the closure of geodesic ball $\mathcal{B}_{o}(1)$ of radius $1$ and cent ered at $o$. Outside of this ball, the Hessian is strictly positive, and choosing $\delta=2$ ensures that (5.3) in Assumption Assumption is satisfied. This implies that choosing $b$ satisfying $0<b<(2(1+\sqrt{2})\sqrt{m-1})^{-1/2}$ takes care of Assumption Assumption. Non-convexity of $\phi$ can now be gleaned from the fact that inside of $\mathcal{B}_{o}(1/\sqrt{2})$ the Hessian is negative. To see this, we look at the Hessian contracted with vector fields $v$ such that $g(v,\nabla\rho)=0$. Then
@@ 5.2. Computational aspects
However, distributions $\phi$ and manifolds $(M,g)$ of practical interest are usually so that the globally Lipschitz assumption is violated. It is even so for Example 5.2(i) of the Riemannian-Gaussian distribution on $\mathcal{P}_{m}$.
@@
In Section 6.2 we provide numerical evidence of how the rejection technique can profitably be used by choosing a large enough closed geodesic ball, as in Assumption Assumption, on $\mathcal{P}_{m}$ using the proposed Riemannian Langevin Monte Carlo algorithm. In particular, no trajectory was rejected for $R=2.7$ in the case of the Riemannian-Gaussian distribution (Section 6.2.1) and $R=2$ for the non-convex potential (Section 6.2.2).
@@
Alternatives to the explicit Euler, that result in bounded moments and hence convergence, have been studied in depth for SDEs in $\mathbb{R}^{q}$ with non-globally Lipschitz coefficients. Examples include implicit methods (and in particular, the implicit Euler scheme) and explicit methods of tamed and balanced type (e.g., Mattingly et al., 2002; Hutzenthaler and Jentzen, 2015; Tretyakov and Zhang, 2013; Milstein and Tretyakov, 2021). Extensions of such methods to the non-compact manifold setting are possible and will be taken up elsewhere.
@@ 6. Numerical Illustrations
We carry out numerical experiments to verify the theoretical results from Section 4 for compact manifolds $M$, and demonstrate utility of the proposed algorithm for both compact and non-compact $M$. Section 6.1 considers sampling using retractions from the von-Mises distribution on the two-dimensional unit sphere $\mathbb{S}^{2}$. A sampling algorithm using extrinsic embedding coordinates was proposed by Wood (1994), while Fisher et al. (1993) proposed an exact method using intrinsic coordinates for dimension $2$. A more recent development in Kurz and Hanebeck (2015) provides an extrinsic algorithm without rejection that is exact for even dimension spheres only.
Section 6.2 considers two distributions on the non-compact manifold of symmetric positive definite matrices: the Riemannian-Gaussian distribution with convex $\phi$, and a distribution with non-convex potential. Versatility of the proposed algorithm to handle both convex and non-convex potential is demonstrated.
We focus on the ensemble-averaging estimator $\hat{\mu}_{\phi,N}(\varphi)$ defined in (4.1). To this end, we simulate a large number $L$ of independent trajectories for a sufficiently long time $T=Nh$, and then verify that it enjoys first order of convergence in terms of time step $h$ by comparing with the exact value of $\mu_{\phi}(\varphi)$ (obtained via numerical integration). Two types of errors are reported: the numerical integration error
@@ 6.1. Retraction-based sampling from the von-Mises Fisher distribution on $\mathbb{S}^{2}$
which is obtained by pulling back the Euclidean metric from $\mathbb{R}^{3}$. The inverse metric tensor assumes the form
@@
The cent ered von-Mises Fisher distribution $\mu_{\phi}$ on $\mathbb{S}^{2}$ is the distribution
@@
so that $\phi((r,\theta)^{\intercal})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda\cos r$; performance of the algorithm for a non-cent ered version can be examined with obvious modifications. It was verified in Example 2.1 that the Langevin diffusion (2.7) with $\phi((r,\theta)^{\intercal})=-\lambda\cos r$ is ergodic, and converges exponentially fast to its invariant measure $\mu_{\phi}$.
@@
Since an exponential map in closed form is not available for this example, we use the retraction based on solving the above ODEs by a single step of size $\sqrt{h}$ of the standard $(1/6,1/3,1/3,1/6)$ 4th-order Runge-Kutta method (RK4) as explained in Section 3.2.
@@
For the testing purposes, we take $\varphi(x)=\sin r$ and compute, for $\lambda=1$, the exact value $\mu_{\phi}(\varphi)=\sqrt{2}\frac{I_{1}(1)\Gamma(3/2)}{I_{1/2}(1)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\approx 0.7554024361}$, where $I_{n}(\lambda)$ is the modified Bessel function of the first kind (Lewis, 2023, Section 4.1). The algorithm was terminated at time $T=5$, sufficient to ensure that the error $|\mathbb{E}(\varphi(X(T)))$ $-\mu_{\phi}(\varphi)|$ (i.e., the error due to proximity of the SDE’s solution to the ergodic limit) is negligibly small. From Table 1 and the black line in Figure 1, we observe the first order convergence of Algorithm 6.1 as expected.
Table 1. Estimation error err and Monte Carlo error MCerr for sampling from a von-Mises Fisher distribution depending on the time step $h$ and number of independent realisations $L$. The parameters are $\lambda=1$, $X_{0}=(r_{0},\theta_{0})^{\intercal}=(\pi/4,\pi/4)^{\intercal}$ and $T=5$.
@@
A further experiment was implemented in Julia using GPU kernels and run on an NVIDIA A100 GPU to compute the second order estimate of $\mu_{\phi}(\varphi)$ using the Talay-Tubaro extrapolation method detailed in Section 4.4. For the initial first-order estimates, the experiment was repeated with the time steps $h=0.1,0.08,0.0625,0.05$ and with parameters $T=10$ and $L=5\times 10^{9}$ ($N_{1}=100,\,N_{2}=125,\,N_{3}=160,\,N_{4}=200$). The respective first-order estimates of $\mu_{\phi}(\varphi)$ were $\hat{\mu}_{\phi,N_{1}}(\varphi)=0.762017\pm 0.000006$ (err $\doteq 0.006615$), $\hat{\mu}_{\phi,N_{2}}=0.760524\pm 0.00006$ (err $\doteq 0.005122$), $\hat{\mu}_{\phi,N_{3}}(\varphi)=0.759317\pm 0.000006$ (err0.003915approaches-limitabsent0.003915$\doteq$ 0.003915≐ 0.003915) and $\hat{\mu}_{\phi,N_{4}}(\varphi)=0.758523\pm 0.000006$ (err $\doteq 0.003121$). Then, using (4.12) and the above data, the three second order estimates of $\mu_{\phi}(\varphi)$ are: $\hat{\mu}_{\phi,imp}(\varphi)=0.754552\pm 0.000041$ (err $\doteq 0.00085$) based on $\hat{\mu}_{\phi,N_{1}}(\varphi)$ and $\hat{\mu}_{\phi,N_{2}}(\varphi)$; $\hat{\mu}_{\phi,imp}(\varphi)=0.755006\pm 0.000037$ (err $\doteq 0.000396$) based on $\hat{\mu}_{\phi,N_{2}}(\varphi)$ and $\hat{\mu}_{\phi,N_{3}}(\varphi)$; and $\hat{\mu}_{\phi,imp}(\varphi)=0.755348\pm 0.000041$ (err $\doteq 0.000054$) based on $\hat{\mu}_{\phi,N_{3}}(\varphi)$ and $\hat{\mu}_{\phi,N_{4}}(\varphi)$. We see that the Talay-Tubaro extrapolation improves the accuracy as expected. The combined computation time of this experiment was about 3000 seconds thanks to an effective use of GPU.